OpenStudy (boat25):

Holly wants to save money for an emergency. Holly invests $600 in an account that pays an interest rate of 7.5%. How many years will it take for the account to reach $7,600? Round your answer to the nearest hundredth.

OpenStudy (welshfella):

Is the interest compounded each year?

OpenStudy (boat25):

Doesn't say but I assume so.

OpenStudy (boat25):

I keep getting stuck on these questions

OpenStudy (aakashtomar):

Well, @welshfella is right, we need to know if it's simple interest or compound interest.

OpenStudy (boat25):

Thats all the question gives me

OpenStudy (boat25):

P = amount of money invested r = interest rate percentage in decimal form t = time in years A = total money in the account at time t A = P (1 + r )^t I use this formula that it gives me and keeps saying incorrect

OpenStudy (aakashtomar):

Then try using \[I = (P \times R \times T) / 100\]

OpenStudy (boat25):

I=(600 * .75 * T)/100 ?